# Informativity Conditions for Multiple Signals: Properties, Experimental Design, and Applications (extended version)

### v3

@@ 4 Experimental Design for CPE Conditions
Recent finite-time experiment design methods (e.g., Venkatasubramanian et al. (2025)) optimize a single exploration trajectory, typically harmonic or multisine inputs, to minimize parameter uncertainty and guarantee a priori performance bounds after identification. In contrast, the goal of our experiment design is not to minimize estimation error for a pre-specified model, but rather to ensure that the data collected across possibly multiple trials satisfy the proposed CPE conditions, which are sufficient for data-driven controller synthesis. This is crucial in settings where a single long PE trajectory cannot be executed due to safety or operational constraints. Moreover, our approach improves the numerical conditioning of the resulting Hankel matrix via weighting, thereby reducing the number of experimental repetitions needed to achieve data informativity. Hence, while targeted exploration focuses on minimizing model uncertainty for robust performance, our design explicitly guarantees the feasibility of direct data-driven control synthesis from multiple finite trajectories, even under limited excitation budgets.
@@ 5 Application in Distributed Adaptive Identification
    $z_{[k,k+l-1]}$ is PE of order $1$ for a fixed constant $l\geq 1$.
@@ 5.1 Adaptive Identifier for a Single Control Sequence Satisfying PE Condition
$\theta(k+1)=\theta(k)+\alpha r(k){\left({{r^{T}}(k)r(k)+\xi I}\right)^{-1}}\varepsilon(k),$ (9)
where $\theta(k)$ is the identification vector at moment $k$, $\alpha>0$ denotes the adaptive gain, $\xi>0$ is an arbitrary positive constant, and $\varepsilon(k)={x(k+1)-{r^{T}}(k)\theta(k)}$.
@@
$\tilde{\theta}(k+1)=(I-\alpha r(k){\left({{r^{T}}(k)r(k)+\xi I}\right)^{-1}}{r^{T}}(k))\tilde{\theta}(k).$ (10)
Since ${r^{T}}(k)r(k)+\xi I$} is an invertible symmetric matrix, (10) possesses the same form as (7). We can use Lemma 3 to establish the exponential convergence condition of the identifier (9), as shown in the following theorem.
@@
    $r_{[k,k+l-1]}$ satisfies PE condition of order $1$ with a fixed constant $l\geq 1$.
Proof: Utilizing the Woodbury Matrix Identity, we obtain $r{({r^{T}}r+\xi I)^{-1}}{r^{T}}=I-{(I+\frac{1}{\zeta}r{r^{T}})^{-1}}<I$. Consequently,
$2I-\alpha r(k){\left({{r^{T}}(k)r(k)+\xi I}\right)^{-1}}{r^{T}}(k)>0$ (11)
@@
Note that for the systems described by (1) or (8), when a set of input-state data can uniquely determine $A$ and $B$, it is typically necessary for the input sequence to constitute a persistently exciting sequence of order $n+1$. If the excitation condition for $u_{[k,k+l-1]}$ holds for all $k\geq 0$ with a fixed constant $l\geq 1$, then $\begin{bmatrix}{{H_{1}}({x_{[k,k+l-1]}})}\\ {{H_{1}}({u_{[k,k+l-1]}})}\end{bmatrix}$ achieves full row rank. Consequently, ${H_{1}}({r_{[k,k+l-1]}})$ also attains full row rank, meeting the excitation requirement outlined in Theorem 3. Therefore, the following corollary ensues.
@@
    $u_{[k,k+l-1]}$ satisfies PE condition of order $n+1$ with a fixed constant $l\geq 1$.
@@ 5.2 Distributed Adaptive Identifiers for Multiple Control Sequences Satisfying CPE Conditions
Using the rank condition under CPE provided in Lemma 1, there is a centralized non-adaptive identification scheme similar to (4). Taking the CCPE condition as an example, one obtains
${H_{1}^{c}({x_{1[1,{T_{0}}]}^{p}})}=\begin{bmatrix}A&B\end{bmatrix}\begin{bmatrix}{H_{1}^{c}({x_{1[0,{T_{0}-1}]}^{p}})}\\ {H_{1}^{c}({u_{1[0,{T_{0}-1}]}^{p}})}\end{bmatrix}$ (13)
If $T_{0}=T_{1}=T_{2}=...=T_{p}$ and $\{u_{i[0,{T_{i}-1}]}\}_{i=1}^{p}$ are CCPE of order $n+1$, then $\begin{bmatrix}{H_{1}^{c}({x_{1[0,{T_{0}-1}]}^{p}})}\\ {H_{1}^{c}({u_{1[0,{T_{0}-1}]}^{p}})}\end{bmatrix}$ has full row rank. Hence, we have
$\begin{bmatrix}A&B\end{bmatrix}={H_{1}^{c}({x_{1[1,{T_{0}}]}^{p}})}{\begin{bmatrix}{H_{1}^{c}({x_{1[0,{T_{0}-1}]}^{p}})}\\ {H_{1}^{c}({u_{1[0,{T_{0}-1}]}^{p}})}\end{bmatrix}^{\dagger}}$, (14)
where $\dagger$} denotes the right inverse.
However, this centralized non-adaptive identification method is complex and poses a higher risk of being vulnerable to covert and stealthy attacks. Additionally, it exhibits poor identification capability in the presence of changing parameters. Therefore, in the following, we delve into the design of distributed adaptive identifiers to tackle these challenges.
@@
${\theta_{i}}(k+1)={\theta_{i}}(k)+$ $\alpha{r_{i}}(k){\left({{r_{i}}^{T}(k){r_{i}}(k)+\xi I}\right)^{-1}}{\varepsilon_{i}}(k)$
@@
where $\theta_{i}(k)$ is the identification vector of $i$-th agent at moment $k$, $\alpha>0$ and $\gamma>0$ are constants to be designed, $\xi>0$ is an arbitrary positive constant, and $\varepsilon_{i}(k)={x_{i}(k+1)-{r_{i}^{T}}(k)\theta_{i}(k)}$.
Let ${{\tilde{\theta}}_{i}}(k)={\theta_{i}}(k)-\theta$ and ${\tilde{\theta}_{c}}(k)=\begin{bmatrix}{\tilde{\theta}_{1}^{T}(k)},{\tilde{\theta}_{2}^{T}(k)},...,{\tilde{\theta}_{N}^{T}(k)}\end{bmatrix}^{T}$ denote the identification error of the $i$-th identifier and the compact form of the errors, respectively. Then, the error dynamics can be expressed as
@@
where $R(k)=\text{diag}(\bar{r}_{1}(k),...,\bar{r}_{N}(k))$ and $\bar{r}_{i}(k)=r_{i}(k){\left({{r_{i}^{T}}(k)r_{i}(k)+\xi I}\right)^{-1}}{r_{i}^{T}}(k)$.
@@
    The inputs $\{u_{i[k,{k+l-1}]}\}_{i=1}^{N}$ are any type of CPE of order $n+1$ for a fixed constant $l\geq 1$.
Proof: By definition, we know that $\alpha R(k)+\gamma{\mathcal{L}}\otimes I$ is a semi-positive definite matrix, so there exists a matrix $\Phi(k)$ such that $\Phi(k)\Phi^{T}(k)=\alpha R(k)+\gamma{\mathcal{L}}\otimes I$. Then, (16) can be written as
$\tilde{\theta}_{c}(k+1)=\left({I-\Phi(k)\Phi^{T}(k)}\right)\tilde{\theta}_{c}(k).$ (17)
We first prove that $\Phi_{[k,k+l-1]}$ is PE of order $1$ for a fixed constant $l\geq 1$ when $\{u_{i[k,{k+l-1}]}\}_{i=1}^{N}$ are MCPE of order $n+1$. The problem can be translated into proving that
$\sum\limits_{t=0}^{l-1}{\Phi(k+t){\Phi^{T}}(k+t)}=l\gamma\mathcal{L}\otimes I+\alpha\sum\limits_{t=0}^{l-1}{R(k+t)}>0.$
@@
${\eta^{T}}{({{\bf{1}}_{(n+m)n}}\otimes{I_{(n+m)n}})^{T}}\sum\limits_{t=0}^{l-1}{R(k+t)}({{\bf{1}}_{(n+m)n}}\otimes{I_{(n+m)n}})\eta$
@@
If $\{u_{i[k,{k+l-1}]}\}_{i=1}^{N}$ satisfy the MCPE condition, it can be readily inferred that $\sum\limits_{t=0}^{l-1}{\sum\limits_{i=1}^{N}{{{\bar{r}}_{i}}(k+t)}}>0$. Thus, (5.2) equals 0 if and only if $\eta$ is the zero vector, implying that $\ker\mathcal{L}\otimes I\cap\ker\sum\limits_{t=0}^{l-1}{R(k+t)}=\{0\}$. Moreover, according to Lemma 1, if $\{u_{i[k,{k+l-1}]}\}_{i=1}^{N}$ satisfy the CCPE or HCPE condition, they must also satisfy the MCPE condition. Therefore, $\Phi_{[k,k+l-1]}$ are PE of order $1$ for all types of CPE of order $n+1$.
Next we prove that $2I-\Phi^{T}(k)\Phi(k)>0$. Let $\alpha R(k)=\phi_{1}(k)\phi_{1}^{T}(k)$ and $\gamma{\mathcal{L}}\otimes I=\phi_{2}(k)\phi_{2}^{T}(k)$, thus we have $\Phi(k)=\begin{bmatrix}{\phi_{1}(k)}&{\phi_{2}(k)}\end{bmatrix}$. Then, the inequality $2I-\Phi^{T}(k)\Phi(k)>0$ is equivalent to
$\begin{bmatrix}{2I-\phi_{1}^{T}(k){\phi_{1}}(k)}&-{\phi_{1}^{T}(k){\phi_{2}}(k)}\\ -{\phi_{2}^{T}(k){\phi_{1}}(k)}&{2I-\phi_{2}^{T}(k){\phi_{2}}(k)}\end{bmatrix}>0.$ (19)
@@
$2I-\phi_{2}^{T}(k){\phi_{2}}(k)-\phi_{2}^{T}(k)M_{2}(k){\phi_{2}}(k)>0,$
where $M_{2}(k)={\phi_{1}}(k){\left({2I-\phi_{1}^{T}(k){\phi_{1}}(k)}\right)^{-1}}\phi_{1}^{T}(k)$. According to the Woodbury Matrix Identity, ${M_{2}}(k)=I-{\left({I+\frac{1}{2}{\phi_{1}}(k)\phi_{1}^{T}(k)}\right)^{-1}}<I$, which implies that
$2I-\phi_{2}^{T}(k){\phi_{2}}(k)-\phi_{2}^{T}(k){M_{2}}(k){\phi_{2}}(k)>2I-2\phi_{2}^{T}(k){\phi_{2}}(k).$
@@ 6.1 Distributed Identification from Multiple Control Inputs
In this subsection, we apply the distributed adaptive identification scheme of Section 5 to the voltage converter systems. Consider a homogeneous multi-agent systems comprising five agents, and the communication network $\mathcal{G}$ is shown in Fig. 2. By discretizing the voltage converter system described in Tan and Hoo (2015) with a sampling time of 0.1 seconds, we derive the state-space representation for each agent as
@@
where $i\in\{1,2,3,4,5\}$. Transforming the system into the form of (12) results in $\theta=\begin{bmatrix}1.0000&{-0.0500}&{0.0125}&{0.0004}&{0.9998}&0.0000\end{bmatrix}^{T}$, where $\theta$ represents the parameters subject to identification.
@@ 6.2 Least-squares Identification from Multiple Control Inputs
In this subsection, we apply the LS identification method to a batch reactor system, previously considered in De Persis and Tesi (2020). The system dynamics are described by
@@ 6.3 State Feedback Controller Design Using Parametric Modeling Approach
In this subsection, we apply the data-driven state feedback design method described in De Persis and Tesi (2020) to the system defined by (6.2) using the three CPE conditions. Specifically, taking the CCPE condition as an example. Utilizing the approach in (De Persis and Tesi, 2020, Th. 3), the design challenge for stabilizing controllers is transformed into an LMI problem for any matirx $Q$:
@@
In the case of the CCPE condition, the collected input-state trajectories are synthesized into $H_{1}^{cum}(\{{u_{i}^{[0,{T_{0}-1}]}\}_{i=1}^{p}})$}, $H_{1}^{cum}(\{{x_{i}^{[0,{T_{0}-1}]}\}_{i=1}^{p}})$}, and $H_{1}^{cum}(\{{x_{i}^{[1,{T_{0}}]}\}_{i=1}^{p}})$}. These matrices are then processed using MATLAB’s LMI toolbox Duan and Yu (2013) to compute the control gain $K_{cum}$ using (22) and (6.3):

### v4

@@ 4 Experimental Design for the CPE Conditions
Recent finite-time experiment design methods (e.g., 10601334) optimize a single exploration trajectory, typically harmonic or multisine inputs, to minimize parameter uncertainty and guarantee a priori performance bounds after identification. In contrast, the goal of our experiment design is not to minimize estimation error for a pre-specified model, but rather to ensure that the data collected across possibly multiple trials satisfy the proposed CPE conditions, which are sufficient for data-driven controller synthesis. This is crucial in settings where a single long PE trajectory cannot be executed due to safety or operational constraints. Moreover, our approach improves the numerical conditioning of the resulting Hankel matrix via weighting, thereby reducing the number of experimental repetitions needed to achieve data informativity. Hence, while targeted exploration focuses on minimizing model uncertainty for robust performance, our design explicitly guarantees the feasibility of direct data-driven control synthesis from multiple finite trajectories, even under limited excitation budgets.
@@ 5 Application in Distributed Adaptive Identification
    $z^{[k,k+l-1]}$ is PE of order $1$ for a fixed constant $l\geq 1$.
@@ 5.1 Adaptive Identifier for a Single Control Sequence Satisfying PE Condition
$\theta(k+1)=\theta(k)+\alpha r(k){\left({{r^{\top}}(k)r(k)+\xi I}\right)^{-1}}\varepsilon(k),$ (9)
where $\theta(k)$ is the identification vector at moment $k$, $\alpha>0$ denotes the adaptive gain, $\xi>0$ is an arbitrary positive constant, and $\varepsilon(k)={x(k+1)-{r^{\top}}(k)\theta(k)}$.
@@
$\tilde{\theta}(k+1)=(I-\alpha r(k){\left({{r^{\top}}(k)r(k)+\xi I}\right)^{-1}}{r^{\top}}(k))\tilde{\theta}(k).$ (10)
Since ${r^{\top}}(k)r(k)+\xi I$} is an invertible symmetric matrix, (10) possesses the same form as (7). We can use Lemma 3 to establish the exponential convergence condition of the identifier (9), as shown in the following theorem.
@@
    $r^{[k,k+l-1]}$ satisfies PE condition of order $1$ with a fixed constant $l\geq 1$.
Proof: Utilizing the Woodbury Matrix Identity, we obtain $r{({r^{\top}}r+\xi I)^{-1}}{r^{\top}}=I-{(I+\frac{1}{\zeta}r{r^{\top}})^{-1}}<I$. Consequently,
$2I-\alpha r(k){\left({{r^{\top}}(k)r(k)+\xi I}\right)^{-1}}{r^{\top}}(k)>0$ (11)
@@
Note that for the systems described by (1) or (8), when a set of input-state data can uniquely determine $A$ and $B$, it is typically necessary for the input sequence to constitute a persistently exciting sequence of order $n+1$. If the excitation condition for $u_{[k,k+l-1]}$ holds for all $k\geq 0$ with a fixed constant $l\geq 1$, then $\begin{bmatrix}{{H_{1}}({x^{[k,k+l-1]}})}\\ {{H_{1}}({u^{[k,k+l-1]}})}\end{bmatrix}$ achieves full row rank. Consequently, ${H_{1}}({r^{[k,k+l-1]}})$ also attains full row rank, meeting the excitation requirement outlined in Theorem 3. Therefore, the following corollary ensues.
@@
    $u^{[k,k+l-1]}$ satisfies PE condition of order $n+1$ with a fixed constant $l\geq 1$.
@@ 5.2 Distributed Adaptive Identifiers for Multiple Control Sequences Satisfying CPE Conditions
${\theta_{i}}(k+1)={\theta_{i}}(k)+$ $\alpha{r_{i}}(k){\left({{r_{i}}^{\top}(k){r_{i}}(k)+\xi I}\right)^{-1}}{\varepsilon_{i}}(k)$
@@
where $\theta_{i}(k)$ is the identification vector of $i$-th agent at moment $k$, $\alpha>0$ and $\gamma>0$ are constants to be designed, $\xi>0$ is an arbitrary positive constant, and $\varepsilon_{i}(k)={x_{i}(k+1)-{r_{i}^{\top}}(k)\theta_{i}(k)}$.
Let ${{\tilde{\theta}}_{i}}(k)={\theta_{i}}(k)-\theta$ and ${\tilde{\theta}_{c}}(k)=\begin{bmatrix}{\tilde{\theta}_{1}^{\top}(k)},{\tilde{\theta}_{2}^{\top}(k)},...,{\tilde{\theta}_{N}^{\top}(k)}\end{bmatrix}^{\top}$ denote the identification error of the $i$-th identifier and the compact form of the errors, respectively. Then, the error dynamics can be expressed as
@@
where $R(k)=\text{diag}(\bar{r}_{1}(k),...,\bar{r}_{N}(k))$ and $\bar{r}_{i}(k)=r_{i}(k){\left({{r_{i}^{\top}}(k)r_{i}(k)+\xi I}\right)^{-1}}{r_{i}^{\top}}(k)$.
@@
    The inputs $\{u_{i}^{[k,{k+l-1}]}\}_{i=1}^{N}$ are any type of CPE of order $n+1$ for a fixed constant $l\geq 1$.
Proof: By definition, we know that $\alpha R(k)+\gamma{\mathcal{L}}\otimes I$ is a semi-positive definite matrix, so there exists a matrix $\Phi(k)$ such that $\Phi(k)\Phi^{\top}(k)=\alpha R(k)+\gamma{\mathcal{L}}\otimes I$. Then, (14) can be written as
$\tilde{\theta}_{c}(k+1)=\left({I-\Phi(k)\Phi^{\top}(k)}\right)\tilde{\theta}_{c}(k).$ (15)
We first prove that $\Phi^{[k,k+l-1]}$ is PE of order $1$ for a fixed constant $l\geq 1$ when $\{u_{i}^{[k,{k+l-1}]}\}_{i=1}^{N}$ are MCPE of order $n+1$. The problem can be translated into proving that
$\sum\limits_{t=0}^{l-1}{\Phi(k+t){\Phi^{\top}}(k+t)}=l\gamma\mathcal{L}\otimes I+\alpha\sum\limits_{t=0}^{l-1}{R(k+t)}>0.$
@@
${\eta^{\top}}{({{\bf{1}}_{(n+m)n}}\otimes{I_{(n+m)n}})^{\top}}\sum\limits_{t=0}^{l-1}{R(k+t)}({{\bf{1}}_{(n+m)n}}\otimes{I_{(n+m)n}})\eta$
@@
If $\{u_{i}^{[k,{k+l-1}]}\}_{i=1}^{N}$ satisfy the MCPE condition, it can be readily inferred that $\sum\limits_{t=0}^{l-1}{\sum\limits_{i=1}^{N}{{{\bar{r}}_{i}}(k+t)}}>0$. Thus, (5.2) equals 0 if and only if $\eta$ is the zero vector, implying that $\ker\mathcal{L}\otimes I\cap\ker\sum\limits_{t=0}^{l-1}{R(k+t)}=\{0\}$. Moreover, according to Lemma 1, if $\{u_{i}^{[k,{k+l-1}]}\}_{i=1}^{N}$ satisfy the CCPE or HCPE condition, they must also satisfy the MCPE condition. Therefore, $\Phi^{[k,k+l-1]}$ are PE of order $1$ for all types of CPE of order $n+1$.
Next we prove that $2I-\Phi^{\top}(k)\Phi(k)>0$. Let $\alpha R(k)=\phi_{1}(k)\phi_{1}^{\top}(k)$ and $\gamma{\mathcal{L}}\otimes I=\phi_{2}(k)\phi_{2}^{\top}(k)$, thus we have $\Phi(k)=\begin{bmatrix}{\phi_{1}(k)}&{\phi_{2}(k)}\end{bmatrix}$. Then, the inequality $2I-\Phi^{\top}(k)\Phi(k)>0$ is equivalent to
$\begin{bmatrix}{2I-\phi_{1}^{\top}(k){\phi_{1}}(k)}&-{\phi_{1}^{\top}(k){\phi_{2}}(k)}\\ -{\phi_{2}^{\top}(k){\phi_{1}}(k)}&{2I-\phi_{2}^{\top}(k){\phi_{2}}(k)}\end{bmatrix}>0.$ (17)
@@
$2I-\phi_{2}^{\top}(k){\phi_{2}}(k)-\phi_{2}^{\top}(k)M_{2}(k){\phi_{2}}(k)>0,$
where $M_{2}(k)={\phi_{1}}(k){\left({2I-\phi_{1}^{\top}(k){\phi_{1}}(k)}\right)^{-1}}\phi_{1}^{\top}(k)$. According to the Woodbury Matrix Identity, ${M_{2}}(k)=I-{\left({I+\frac{1}{2}{\phi_{1}}(k)\phi_{1}^{\top}(k)}\right)^{-1}}<I$, which implies that
$2I-\phi_{2}^{\top}(k){\phi_{2}}(k)-\phi_{2}^{\top}(k){M_{2}}(k){\phi_{2}}(k)>2I-2\phi_{2}^{\top}(k){\phi_{2}}(k).$
@@ 6.1 Distributed Identification from Multiple Control Inputs
In this subsection, we apply the distributed adaptive identification scheme of Section 5 to the voltage converter systems. Consider a homogeneous multi-agent systems comprising five agents, and the communication network $\mathcal{G}$ is shown in Fig. 2. By discretizing the voltage converter system described in 7409511 with a sampling time of 0.1 seconds, we derive the state-space representation for each agent as
@@
where $i\in\{1,2,3,4,5\}$. Transforming the system into the form of (12) results in $\theta=\begin{bmatrix}1.0000&{-0.0500}&{0.0125}&{0.0004}&{0.9998}&0.0000\end{bmatrix}^{\top}$, where $\theta$ represents the parameters subject to identification.
@@ 6.2 Least-squares Identification from Multiple Control Inputs
In this subsection, we apply the LS identification method to a batch reactor system, previously considered in 8933093. The system dynamics are described by
@@ 6.3 State Feedback Controller Design Using Parametric Modeling Approach
In this subsection, we apply the data-driven state feedback design method described in 8933093 to the system defined by (6.2) using the three CPE conditions. Specifically, taking the CCPE condition as an example. Utilizing the approach in (8933093, Th. 3), the design challenge for stabilizing controllers is transformed into an LMI problem for any matirx $Q$:
@@
In the case of the CCPE condition, the collected input-state trajectories are synthesized into $H_{1}^{cum}(\{{u_{i}^{[0,{T_{0}-1}]}\}_{i=1}^{p}})$}, $H_{1}^{cum}(\{{x_{i}^{[0,{T_{0}-1}]}\}_{i=1}^{p}})$}, and $H_{1}^{cum}(\{{x_{i}^{[1,{T_{0}}]}\}_{i=1}^{p}})$}. These matrices are then processed using MATLAB’s LMI toolbox duan2013lmis to compute the control gain $K_{cum}$ using (20) and (6.3):